1. a). Write down the definition (ε-δ language) of limx→x0 f(x) = L. b). Show that limx→+∞ cos x does not exist.


1
Expert's answer
2021-05-06T16:18:02-0400

a). Let us write down the definition ("\u03b5-\u03b4" language) of "\lim\limits_{x→x_0} f(x) = L":

"\lim\limits_{x→x_0} f(x) = L" if and only if for any "\varepsilon>0" there exists "\delta>0" such that for any "x\ne x_0" if "|x-x_0|<\delta" then "|f(x)-L|<\varepsilon."


b). Let us show that "\lim\limits_{x→+\infty} \cos x"  does not exist. Suppose contrary that such limit exists and equals to "L". Then for any "\varepsilon>0" there exists "\Delta>0" such that for any "x>\Delta" we have that "|f(x)-L|<\varepsilon."

Let "\varepsilon =1". Then for any even integer number "n>\Delta" we have that "n\pi>\Delta" and "n\pi+\pi>\Delta". Since the period of the function "y=\cos x" is "2\pi", we conclude that "\cos n\pi =\cos 0 =1" and "\cos (n\pi+\pi) =\cos \pi =-1". By definition of limit we have that "|1-L|=|\cos n\pi -L|<1" and "|-1-L|=|\cos (n\pi+\pi) -L|<1". The latter inequalities are equivalent to "0<L<2" and "-2<L<0," that is impossible. This contradiction proves that the limit does not exist.